A chemist has some 12% hydogen peroxide solution and some 9% hydogen peroxide solution. how many milliters of each should be mixed to make a 450-milliter solution that is 10% hydogen peroxide?
Answer by checkley77(12569)

.12X+.09(450-X)=.10*450
.12X+40.5-.09X=45
.03X=45-40.5
.03X=4.5
X=4.5/.03
X=150 ML. OF 12% HYDROGEN PEROXIDE.
450-150=300 ML OF 9% HYROGEN PEROXIDE.
PROOF:
.12*150+.09*300=.10*450
18+27=45
45=45